1. An unknown wire experiences an electric field of 10 V/m, and has a current density of 5.0 ✕ 107 A/m2 . What is the resistivity of the wire?


1
Expert's answer
2022-01-24T08:44:46-0500

Given:

"E=\rm 10\: V/m"

"j=\rm 5.0*10^7\: A/m^2"


The Ohmn law says

"j=\sigma E=\frac{1}{\rho}E"

Hence, the resistivity of a wire

"\rho=\frac{E}{j}=\frac{10}{5.0*10^7}=2.0*10^{-7}\:\rm \Omega\cdot m"